# Adaptive Real-Time Scheduling Algorithms for Embedded Systems

### v1

@@ I Introduction
The embedded systems that are used in modern times, from automated vehicles to smart medical equipment, are increasingly dynamic environments where the workloads, availability of resources, and external conditions vary dynamically. Conventional real-time scheduling algorithms, such as Rate-Monotonic (RM) and Earliest Deadline First (EDF), make an additional assumption that the set of tasks is known and that the execution times are fixed, where they cannot cope with uncertainty at runtime without making timing guarantees [datta2022adaptive].
In autonomous systems, such as next-generation robots, adaptive path planning algorithms like hybrid PSO-APF approaches [benmachiche2025adaptive] demonstrate the need for real-time adaptation in dynamic environments with static and dynamic obstacles. Similarly, hybrid BFO/PSO algorithms for mobile robot navigation [makhlouf2024enhanced] showcase how bio-inspired optimization can enhance real-time decision-making in embedded systems.
To bridge this gap, adaptive real-time scheduling has gained significant traction, enabling systems to modify priorities, execution budgets, or resource allocations at runtime based on feedback, prediction, or environmental sensing. Among the promising directions are feedback control scheduling (FCS) frameworks that use control-theoretic loops to maintain performance under load variations [lu2002feedback], and machine learning–enhanced predictive schedulers that leverage historical data to anticipate task behavior and proactively adjust schedules [Gracias]. The integration of adaptation based on machine learning into on-board safety-critical systems is of particular importance. Furthermore, the detection of real-time anomalies of onboard aircraft devices requires that the timing and accuracy of detection be optimized for timing and accuracy while respecting the strictest limits of energy consumption and latency constraints [ben].
Recent research highlights this development, for instance Subramaniyan et al. proposed FC-GPU [subramaniyan2025fc], a real-time embedded feedback control scheme that dynamically controls the allocation of GPU resources to soft deadline-based real-time edge AI applications, an essential feature of edge AI applications. In the meantime, Goksoy suggested a runtime monitoring system of ML-based schedulers that indicates a change of distribution during the execution of the task and allows a safe adjustment of safety-critical environments [goksoy2024runtime]. In another study [zhao2024task], Zhao revealed a task-degradation-aware adaptive scheduler that is effective in managing the dynamically applications and maintaining the real-time restrictions.
@@ II-A Feedback-Based Scheduling
One of the foundational structures in this area was presented by Lu et al. [lu2002feedback] with the concept of Feedback Control Scheduling (FCS), the application of classical control theory to real-time systems. Their work enabled the systematic modeling of computing systems as dynamical systems and the development of controllers to control ratios of CPU utilization and deadline misses. This paper showed that analytically tuned feedback controllers can offer high levels of transient and steady-state performance even in cases where task execution times change by large factors at runtime.
In another research paper by Subramaniyan and Wang [subramaniyan2025fc], FC-GPU (Feedback Control GPU Scheduling) is introduced, the first feedback-based scheduling model specifically developed to fit GPUs in embedded and real-time systems. They use a multi-input, multi-output (MIMO) system to model the contention of resources in GPUs to extend feedback control scheduling to the heterogeneous architecture of a GPU. A MIMO controller is a dynamically adjusted task invocation rate controller, which adjusts the rate in response to measured response times. According to the experimental findings on NVIDIA RTX 3090 and AMD MI-100 GPUs, both show considerable enhancements in real-time and high-endurance performance under fluctuations in workloads during runtime.
Pan and Wei proposed a real-time workflow feedback scheduling system in cloud and distributed systems using deep reinforcement learning. They apply a regularized Deep Q-Network (R-DQN) that assigns workflow tasks to virtual machines depending on the condition of the system. This architecture proves to be more adaptive, scalable, and robust than conventional heuristic and fixed schedulers, which essentially serve as a smart feedback controller that learns the best scheduling policies online [pan2024deep].
Scheduling of industrial processes has also been done successfully using feedback principles. He et al. [he2025closed] suggested a closed-loop gasoline blending scheduling scheme that unites real-time optimization and slack-based feedback. Their approach adds slack variables to the constraints of the processes and feeds actual deviations back into the scheduling model and re-optimizes production plans while accounting for actual deviations. Such a strategy enhances the similarity between theoretical scheduling and the real behavior of operations in a highly dynamic environment.
@@ II-B Symbiotic / Task-Coupling Scheduling
Posluns and Jeffrey [posluns2025symbiotic], in their article about Symbiotic Task Scheduling and Data Prefetching, introduce a modern architectural view of the concept of symbiotic scheduling. They present the Task-Seeded Prefetcher (TSP) and Memory Response Task Scheduler (MRS), a joint hardware-software system, which allows task scheduling and memory subsystem behavior to cooperate. TSP acquires task-specific data access patterns and prefetches memory at the granularity of short-lived tasks, whereas MRS takes short-lived tasks as the source of prefetch status decisions. This symbiotic association achieves significant reductions in exposed DRAM latency and substantial performance gains in large-scale manycore processors.
Symbiotic scheduling was first investigated in the context of Simultaneous Multithreading (SMT) architectures in the form of the Symbiotic Job Scheduling (SOS) scheme by Snavely and Tullsen [articlsnavelye]. SOS is a dynamic approach that finds workloads that run effectively as job co-scheduling combinations by sampling different combinations that may or may not work effectively. Through intelligent pairing of threads by using online sampling and a symbiosis-based scheduling phase, SOS proved to achieve better system throughput and lower response time by exploiting complementary usage of resources by threads.
Symbiotic concepts have been applied to real-time IoT workloads in distributed and edge-cloud environments. A semi-dynamic and real-time multiplexing algorithm of task scheduling in cloud-fog IoT systems was proposed by Abohamama et al. [abohamama2022real], where implicit task coupling is observed by optimization by means of permutation based on a modified genetic algorithm. The algorithm clusters and ranks tasks to maximize execution locality and resource fit between the fog and the cloud nodes, resulting in significant gains in makespan, latency, and failure rate for delay-sensitive IoT applications.
Recently, Kwon et al. [kwon2025real] applied the concept of task-coupling to Industrial IoT (IIoT)-based flexible manufacturing systems that involve human-machine interaction. Their model considers operator-induced variation of tasks as explicit coupling events and pre-calculates executable joint resource strategies that combine CPU frequency scaling, memory allocation, and edge/cloud offloading choices. At runtime, the scheduler alternates between these coupled plans on the basis of perceived interactions, providing strict adherence to deadlines and optimizing energy use and system responsiveness.
@@ II-C Predictive / ML-Based Scheduling
Pan and Wei [pan2024deep] designed a representative cloud-based model and suggested a deep learning–based reinforcement learning scheduling system in real-time workflow applications. Their architecture uses a regularized Deep Q-Network (R-DQN), which dynamically plans workflow tasks to virtual machine instances depending on the perceived system state. The approach performs well in optimizing workflows for workload uncertainty and increased resource usage in a heterogeneous workflow structure because it learns the best scheduling policies online, thus making it highly scalable.
In another study, Kesavan and his colleagues [kesavan2025reinforcement] proposed a model named Secure Edge-Enabled Multi-Task Scheduling (SEE-MTS) in edge- and IoE-based settings that combines reinforcement learning and security- and energy-conscious scheduling. Their system integrates edge computing and encrypted task management, dynamic key generation and verification mechanisms, and reinforcement learning to reduce task completion time and energy utilization. The framework is energy-efficient and offers strong security and assurance, which demonstrates the appropriateness of ML-based scheduling for safety-critical and resource-constrained distributed systems.
Conceptually and in terms of systems-level, Guo [guo2025machine] explored the combination of neural networks and machine learning methods with real-time scheduling. This paper points out how neurodynamic systems and reinforcement learning can be used to solve constrained optimization problems subject to time constraints in real time, and the architectural and safety issues of determinism, hardware accelerators, and certification in safety-critical workloads. The research points out the significance of co-designing learning algorithms along with real-time system guarantees.
Related AI approaches in other domains include intrusion detection systems using deep learning [sedraoui2024intrusion] and voice recognition platforms using convolutional neural networks [benmachiche2022development], which demonstrate the potential of ML techniques for real-time embedded applications. Furthermore, optimization techniques like bacterial foraging optimization (BFO) applied to Hidden Markov Models for speech recognition [benmachiche2020optimization, benmachiche2019optimization] illustrate how adaptive algorithms can optimize system parameters in real-time applications.
Similarly, Gracias and Brooklyn [Gracias] discussed AI and ML applications in predictive scheduling within real-time operating systems (RTOS). Their work uses supervised learning models to estimate the execution time of tasks, reinforcement learning to produce adaptive scheduling policies, and context-aware prioritization of tasks using neural networks. Their findings depict better latency, predictability, and throughput than conventional heuristic schedulers, particularly when dealing with extremely variable and unpredictable workloads.
@@ II-D Hybrid / DVFS-Integrated Scheduling
The researchers in [li2024fidrl] implemented the Flexible Invocation-Based Deep Reinforcement Learning (FiDRL) framework for DVFS scheduling in embedded systems. FiDRL enhances regular DRL methods by embedding the invocation timing of the agent within its action space, which allows the agent to self-invoke judiciously to minimize the overall energy of the system, accounting for the energy overhead of the DRL agent. FiDRL allows inter- and intra-task DVFS scheduling and also uses a hybrid on/off-chip algorithm to train and deploy on an embedded platform, achieving a 55.1% decrease in the energy overhead of agent invocation.
In another study [li2025real], the authors proposed a real-time data-driven hybrid synchronization framework to address heterogeneous demand–capacity synchronization (HDCS) in flexible manufacturing systems. Their approach integrates hierarchical real-time data feedback with a ticket-enabled queuing mechanism (GiMS) to ensure seamless coordination between planning, scheduling, and execution. The framework combines global optimization models with local adaptive control mechanisms, enabling rapid reactions to disturbances while preserving overall system optimality.
For flexible and multiskilled manufacturing systems with unpredictable demand, Xinyi Li et al. [li2025energy] defined a real-time data-driven hybrid framework integrating planning, scheduling, and execution (PSE). A stratified real-time feedback data system with hybrid optimization is included in the framework to achieve efficiency and responsiveness of local and global performance simultaneously through the use of human and machine cooperation. Experimental results show cost-efficiency, timeliness, and good resource management, underscoring how hybrid scheduling with real-time data feedback optimizes energy, time, and operational performance in Industry 5.0.

### v2

@@ I Introduction
The embedded systems that are used in modern times, from automated vehicles to smart medical equipment, are increasingly dynamic environments where the workloads, availability of resources, and external conditions vary dynamically. Conventional real-time scheduling algorithms, such as Rate-Monotonic (RM) and Earliest Deadline First (EDF), make an additional assumption that the set of tasks is known and that the execution times are fixed, where they cannot cope with uncertainty at runtime without making timing guarantees [7].
In autonomous systems, such as next-generation robots, adaptive path planning algorithms like hybrid PSO-APF approaches [2] demonstrate the need for real-time adaptation in dynamic environments with static and dynamic obstacles. Similarly, hybrid BFO/PSO algorithms for mobile robot navigation [18] showcase how bio-inspired optimization can enhance real-time decision-making in embedded systems.
To bridge this gap, adaptive real-time scheduling has gained significant traction, enabling systems to modify priorities, execution budgets, or resource allocations at runtime based on feedback, prediction, or environmental sensing. Among the promising directions are feedback control scheduling (FCS) frameworks that use control-theoretic loops to maintain performance under load variations [17], and machine learning–enhanced predictive schedulers that leverage historical data to anticipate task behavior and proactively adjust schedules [9]. The integration of adaptation based on machine learning into on-board safety-critical systems is of particular importance. Furthermore, the detection of real-time anomalies of onboard aircraft devices requires that the timing and accuracy of detection be optimized for timing and accuracy while respecting the strictest limits of energy consumption and latency constraints [6].
Recent research highlights this development, for instance Subramaniyan et al. proposed FC-GPU [23], a real-time embedded feedback control scheme that dynamically controls the allocation of GPU resources to soft deadline-based real-time edge AI applications, an essential feature of edge AI applications. In the meantime, Goksoy suggested a runtime monitoring system of ML-based schedulers that indicates a change of distribution during the execution of the task and allows a safe adjustment of safety-critical environments [8]. In another study [24], Zhao revealed a task-degradation-aware adaptive scheduler that is effective in managing the dynamically applications and maintaining the real-time restrictions.
@@ II-A Feedback-Based Scheduling
One of the foundational structures in this area was presented by Lu et al. [17] with the concept of Feedback Control Scheduling (FCS), the application of classical control theory to real-time systems. Their work enabled the systematic modeling of computing systems as dynamical systems and the development of controllers to control ratios of CPU utilization and deadline misses. This paper showed that analytically tuned feedback controllers can offer high levels of transient and steady-state performance even in cases where task execution times change by large factors at runtime.
In another research paper by Subramaniyan and Wang [23], FC-GPU (Feedback Control GPU Scheduling) is introduced, the first feedback-based scheduling model specifically developed to fit GPUs in embedded and real-time systems. They use a multi-input, multi-output (MIMO) system to model the contention of resources in GPUs to extend feedback control scheduling to the heterogeneous architecture of a GPU. A MIMO controller is a dynamically adjusted task invocation rate controller, which adjusts the rate in response to measured response times. According to the experimental findings on NVIDIA RTX 3090 and AMD MI-100 GPUs, both show considerable enhancements in real-time and high-endurance performance under fluctuations in workloads during runtime.
Pan and Wei proposed a real-time workflow feedback scheduling system in cloud and distributed systems using deep reinforcement learning. They apply a regularized Deep Q-Network (R-DQN) that assigns workflow tasks to virtual machines depending on the condition of the system. This architecture proves to be more adaptive, scalable, and robust than conventional heuristic and fixed schedulers, which essentially serve as a smart feedback controller that learns the best scheduling policies online [19].
Scheduling of industrial processes has also been done successfully using feedback principles. He et al. [11] suggested a closed-loop gasoline blending scheduling scheme that unites real-time optimization and slack-based feedback. Their approach adds slack variables to the constraints of the processes and feeds actual deviations back into the scheduling model and re-optimizes production plans while accounting for actual deviations. Such a strategy enhances the similarity between theoretical scheduling and the real behavior of operations in a highly dynamic environment.
@@ II-B Symbiotic / Task-Coupling Scheduling
Posluns and Jeffrey [20], in their article about Symbiotic Task Scheduling and Data Prefetching, introduce a modern architectural view of the concept of symbiotic scheduling. They present the Task-Seeded Prefetcher (TSP) and Memory Response Task Scheduler (MRS), a joint hardware-software system, which allows task scheduling and memory subsystem behavior to cooperate. TSP acquires task-specific data access patterns and prefetches memory at the granularity of short-lived tasks, whereas MRS takes short-lived tasks as the source of prefetch status decisions. This symbiotic association achieves significant reductions in exposed DRAM latency and substantial performance gains in large-scale manycore processors.
Symbiotic scheduling was first investigated in the context of Simultaneous Multithreading (SMT) architectures in the form of the Symbiotic Job Scheduling (SOS) scheme by Snavely and Tullsen [22]. SOS is a dynamic approach that finds workloads that run effectively as job co-scheduling combinations by sampling different combinations that may or may not work effectively. Through intelligent pairing of threads by using online sampling and a symbiosis-based scheduling phase, SOS proved to achieve better system throughput and lower response time by exploiting complementary usage of resources by threads.
Symbiotic concepts have been applied to real-time IoT workloads in distributed and edge-cloud environments. A semi-dynamic and real-time multiplexing algorithm of task scheduling in cloud-fog IoT systems was proposed by Abohamama et al. [1], where implicit task coupling is observed by optimization by means of permutation based on a modified genetic algorithm. The algorithm clusters and ranks tasks to maximize execution locality and resource fit between the fog and the cloud nodes, resulting in significant gains in makespan, latency, and failure rate for delay-sensitive IoT applications.
Recently, Kwon et al. [13] applied the concept of task-coupling to Industrial IoT (IIoT)-based flexible manufacturing systems that involve human-machine interaction. Their model considers operator-induced variation of tasks as explicit coupling events and pre-calculates executable joint resource strategies that combine CPU frequency scaling, memory allocation, and edge/cloud offloading choices. At runtime, the scheduler alternates between these coupled plans on the basis of perceived interactions, providing strict adherence to deadlines and optimizing energy use and system responsiveness.
@@ II-C Predictive / ML-Based Scheduling
Pan and Wei [19] designed a representative cloud-based model and suggested a deep learning–based reinforcement learning scheduling system in real-time workflow applications. Their architecture uses a regularized Deep Q-Network (R-DQN), which dynamically plans workflow tasks to virtual machine instances depending on the perceived system state. The approach performs well in optimizing workflows for workload uncertainty and increased resource usage in a heterogeneous workflow structure because it learns the best scheduling policies online, thus making it highly scalable.
In another study, Kesavan and his colleagues [12] proposed a model named Secure Edge-Enabled Multi-Task Scheduling (SEE-MTS) in edge- and IoE-based settings that combines reinforcement learning and security- and energy-conscious scheduling. Their system integrates edge computing and encrypted task management, dynamic key generation and verification mechanisms, and reinforcement learning to reduce task completion time and energy utilization. The framework is energy-efficient and offers strong security and assurance, which demonstrates the appropriateness of ML-based scheduling for safety-critical and resource-constrained distributed systems.
Conceptually and in terms of systems-level, Guo [10] explored the combination of neural networks and machine learning methods with real-time scheduling. This paper points out how neurodynamic systems and reinforcement learning can be used to solve constrained optimization problems subject to time constraints in real time, and the architectural and safety issues of determinism, hardware accelerators, and certification in safety-critical workloads. The research points out the significance of co-designing learning algorithms along with real-time system guarantees.
Related AI approaches in other domains include intrusion detection systems using deep learning [21] and voice recognition platforms using convolutional neural networks [3], which demonstrate the potential of ML techniques for real-time embedded applications. Furthermore, optimization techniques like bacterial foraging optimization (BFO) applied to Hidden Markov Models for speech recognition [4, 5] illustrate how adaptive algorithms can optimize system parameters in real-time applications.
Similarly, Gracias and Brooklyn [9] discussed AI and ML applications in predictive scheduling within real-time operating systems (RTOS). Their work uses supervised learning models to estimate the execution time of tasks, reinforcement learning to produce adaptive scheduling policies, and context-aware prioritization of tasks using neural networks. Their findings depict better latency, predictability, and throughput than conventional heuristic schedulers, particularly when dealing with extremely variable and unpredictable workloads.
@@ II-D Hybrid / DVFS-Integrated Scheduling
The researchers in [14] implemented the Flexible Invocation-Based Deep Reinforcement Learning (FiDRL) framework for DVFS scheduling in embedded systems. FiDRL enhances regular DRL methods by embedding the invocation timing of the agent within its action space, which allows the agent to self-invoke judiciously to minimize the overall energy of the system, accounting for the energy overhead of the DRL agent. FiDRL allows inter- and intra-task DVFS scheduling and also uses a hybrid on/off-chip algorithm to train and deploy on an embedded platform, achieving a 55.1% decrease in the energy overhead of agent invocation.
In another study [15], the authors proposed a real-time data-driven hybrid synchronization framework to address heterogeneous demand–capacity synchronization (HDCS) in flexible manufacturing systems. Their approach integrates hierarchical real-time data feedback with a ticket-enabled queuing mechanism (GiMS) to ensure seamless coordination between planning, scheduling, and execution. The framework combines global optimization models with local adaptive control mechanisms, enabling rapid reactions to disturbances while preserving overall system optimality.
For flexible and multiskilled manufacturing systems with unpredictable demand, Xinyi Li et al. [16] defined a real-time data-driven hybrid framework integrating planning, scheduling, and execution (PSE). A stratified real-time feedback data system with hybrid optimization is included in the framework to achieve efficiency and responsiveness of local and global performance simultaneously through the use of human and machine cooperation. Experimental results show cost-efficiency, timeliness, and good resource management, underscoring how hybrid scheduling with real-time data feedback optimizes energy, time, and operational performance in Industry 5.0.